# DLoRA-TrOCR: Mixed Text Mode Optical Character Recognition Based On Transformer

### v1

@@ Abstract
With the continuous development of OCR technology and the expansion of application fields, text recognition in complex scenes has become a key challenge.Factors such as multiple fonts, mixed scenes and complex layouts seriously affect the recognition accuracy of traditional OCR models. Although OCR models based on deep learning have performed well in specific fields or similar data sets in recent years, the generalization ability and robustness of the model are still a big challenge when facing complex environments with multiple scenes. Furthermore, training an OCR model from scratch or fine-tuning all parameters is very demanding on computing resources and inference time, which limits the flexibility of its application. This study focuses on a fundamental aspect of mixed text recognition in response to the challenges mentioned above, which involves effectively fine-tuning the pre-trained basic OCR model to demonstrate exceptional performance across various downstream tasks.To this end, we propose a parameter-efficient hybrid text recognition method based on pre-trained OCR Transformer, namely DLoRA-TrOCR. This method embeds DoRA into the image encoder and LoRA into the internal structure of the text decoder, enabling efficient parameter fine-tuning for downstream tasks.Experimental results show that compared to similar parameter adjustment methods, our model DLoRA-TrOCR has the smallest number of parameters and performs better. It can achieve state-of-the-art performance on complex scene data sets involving simultaneous recognition of mixed handwritten, printed and street view texts.
@@ 1 Introduction
Figure 1: Optical Character Recognition: Text Detection and Text Recognition
@@
Our research aims to extend the application of pre-training and fine-tuning to the OCR domain. To construct a versatile multi-scene text recognition model, we built a dataset containing a mixture of handwritten, printed, and scene text. We then leveraged the pre-trained Transformer OCR model TrOCR, utilizing its weights from training on synthetic handwritten and printed text, and fine-tuned it on our mixed dataset. Introducing parameter-efficient fine-tuning (PEFT) techniques, we adopted the DoRA and LoRA methods for the encoder and decoder, respectively. This significantly reduced the trainable parameters while effectively improving performance on the mixed dataset. Experiments demonstrated outstanding performance in handling complex scenarios with mixed handwritten, printed, and scene text using only a small number of trainable parameters.
@@ 2 Related Work
The SRN proposed by Yu et al[33], employs Transformer-encoder structure followed by the convolutional feature network to enhance visual features. In terms of decoder, the SRACN text detection algorithm proposed by Yang et al[31], uses Transformer-decoder to replace LSTM module, which once again verifies the efficiency and accuracy advantages of parallel training. The NRTR proposed by Sheng et al[25]. uses a complete Transformer architecture to encode and decode input images, and only uses a few simple convolutional layers for deep feature extraction to verify the effectiveness of the Transformer structure in text recognition. The TrOCR model proposed by Li et al. applies the pre-trained Transformer encoder-decoder framework to the OCR task. Unlike previous methods, TrOCR does not use CNN as the backbone network for image understanding, but directly uses the pre-trained image Transformer model as the encoder and the pre-trained text Transformer model as the decoder.
@@
The foundation model[36]has strong generalization ability, and more and more research attempts to use transfer learning to apply it to tasks in specific scenarios. However, the scale of modern deep learning and the requirements for computing scale make traditional transfer learning strategies gradually ineffective, and the need for efficient parameter adjustment becomes increasingly urgent[4]. Parameter-Efficient Fine-Tuning (PEFT)[19] aims to reduce the high cost of fine-tuning large models. This type of method achieves adaptation to downstream tasks by adjusting a relatively small subset of parameters, rather than adjusting all parameters of the model as in traditional methods. This strategy can significantly reduce computational cost and storage requirements compared to the total parameter amount.
Existing PEFT methods can be mainly divided into two categories: addition-based methods and reparameterization-based methods.Our research focuses on the second category of heavy-parameter-based methods, which are implemented by converting the adaptive parameters in the optimization process into a parametrically efficient form. This method was originally applied to natural language models. The representative of this type of method is Low-Rank Adaptation(LoRA)[8], which assumes that changes in weights during model adjustment have low intrinsic rank. Based on this assumption, LoRA proposes a low-order decomposition optimization approximation method for the changes in the original weight matrix in the self-attention module. LoRA can match the fine-tuned performance of the GLUE benchmark while adding almost no additional inference overhead. Subsequently, Liu et al. introduced the idea of weight normalization based on LoRA, decomposed the pre-training weight into two components, amplitude and direction, and used LoRA to update the direction component. This method, called DoRA (Weight-Decomposed Low-Rank Adaptation)[32], enhances the learning capabilities and training stability of LoRA on some tasks without introducing any additional inference overhead.
Figure 2: Transformer-based architecture, where the encoder-decoder model is a pre-trained image transformer as the encoder and a pre-trained text transformer as the decoder.Then, this model is pre-trained in two stages on a synthetic dataset of millions of handwritten and printed texts
@@ 3.1 Transformer-based OCR
Figure 3: Framework diagram of LoRA and DoRA methods.It is intuitive to see that after DoRA decomposes the pre-trained weights into directions and magnitudes, it updates the increments using LoRA in directions before computing with magnitudes
@@ 3.2 Parameter-efficient Fine-tuning
where $B\in\mathbb{R}^{m\times r},A\in\mathbb{R}^{r\times n}$ are two small matrices, called LoRA weights, which are parameters that need to be learned during the fine-tuning process. $r$ is a hyperparameter that controls the rank of LoRA weight. When $r\ll\min(m,n)$, the number of parameters required by the LoRA weight will be much smaller than the original model weight matrix. By adding this low-rank update, LoRA enables efficient fine-tuning of pre-trained models. During the inference phase, the fine-tuned LoRA weights are combined with the original model weights to obtain the final inference weights. Since only two small matrices are stored, LoRA greatly reduces the additional parameter storage space required for fine-tuning. In our text decoder, we apply LoRA to both the query, key, and value projection matrices of the self-attention module, as well as the two linear layers of the feedforward network. This parameter-efficient fine-tuning method allows us to significantly reduce the number of parameters required to fine-tune the text decoder while maintaining performance. Specifically, LoRA’s weight matrix is used in Transformer block to reshape the multi-head attention mechanism, the incremental update of which can be expressed as
$Q=(W_{q}+B_{q}A_{q})\mathbf{x}$
$K=(W_{k}+B_{k}A_{k})\mathbf{x}$
$V=(W_{v}+B_{v}A_{v})\mathbf{x}.$
@@
$Q=(W_{q}+\|W_{q}\|_{c}\frac{W_{q}+B_{q}A_{q}}{\|W_{q}+B_{q}A_{q}\|_{c}})\mathbf{x}$
$K=(W_{k}+\|W_{k}\|_{c}\frac{W_{k}+B_{k}A_{k}}{\|W_{k}+B_{k}A_{k}\|_{c}})\mathbf{x}$
$V=(W_{v}+\|W_{v}\|_{c}\frac{W_{v}+B_{v}A_{v}}{\|W_{v}+B_{v}A_{v}\|_{c}})\mathbf{x}.$
@@
Table 1: Optimization of computational resources and model parameters using LoRA and DoRA methods, which load the pre-trained weights of the model using fp16, can actually be loaded with less precision in practice, which means that a larger batch size can be used on the same memory size, resulting in faster training.
Model PEFT Method GPU Memory Cost Tranable Params Total Params # Params
@@ 4.1 Experiment Setup
From the IAM dataset[20], which contains handwritten English text and is one of the most commonly used datasets in the field of handwritten text recognition. We use the partitioning of the Aachen dataset: the training set contains 6,482 lines of text in 747 forms, the validation set contains 966 lines of text in 115 forms, and the test set contains 2,915 lines of text in 336 forms.
@@
From the task2 of SROIE dataset[9], which focuses on text recognition in receipt images. The SROIE training set and test set contain 626 and 361 receipt images respectively. After cropping and removing duplicate images, the training set contains 10,682 lines of text, and the test set contains 6,897 lines of text.
@@
From multiple public benchmarks, including IIIT5K-Words (IIIT5K)[17], Street View Text(SVT)[29], ICDAR 2013 (IC13)[10], ICDAR 2015(IC15)[11], Street View Text-Perspective (SVTP)[18], and CUTE80 (CUTE)[24]. These datasets contain various scene text images with noise such as blur, occlusion, or low resolution. The training set contains 7573 lines of text, and the test set contains 11435 lines of text. The specific data set situation is shown in the Table 2. To achieve better generalization on multi-scenario tasks, we adopt hybrid measures on the dataset. Compared with the circular sampling method [27]that may lead to errors in uneven data distribution, we adopt a direct and balanced data integration strategy to directly mix the data sets of the above-mentioned multiple related tasks, eliminating the possibility of a single task dominating the training process. nature, promoting the comprehensive learning of the model.
@@
The datasets used in our subsequent comparative experiments, combination experiments, and ablation experiments are all trained and tested on the above-mentioned handwritten text dataset, printed text dataset, scene text dataset, and our mixed text dataset.
@@
We considered two evaluation metrics that are widely used in OCR systems, Character Error Rate(CER) and F1 score. CER is designed to be evaluated at the character level, which is case-sensitive, while the F1 score is evaluated at the word level, which is case-insensitive. CER is defined as the edit distance between the predicted text and the ground truth. CER can be calculated as:
@@
WER and CER are similar benchmark scores. The main difference is that the smallest unit of text in the calculation method is changed from characters to words, so that WAR can be calculated by $1-WER$. We mainly use it in the evaluation of the STR data set.
@@
In the model configuration, we use the official pre-training weights of the TrOCR-Base model, and use the AdamW optimizer to perform training for 20 epochs and a batch size of 16. The initial learning rate of full-parameter Fine-Tuning is 1e-5, and the initial learning rate of PEFT method is 5e-5. Our comparative model experiments are all based on PyTorch, and all experiments are performed on an Nvidia RTX 4090 GPU with 24 GB of memory. In order to accurately evaluate the model performance and ensure the rationality of the training, validation and testing processes, we divided the mixed data set into a training set and a validation set according to the classic 9:1 ratio. In addition, we also retain an independent test set before mixing to ensure that the evaluation of the final performance of the model is completely independent of the training process.
@@ 4.2 Experiment Result
The evaluation covers the above-mentioned benchmark data sets, and uses F1 and CER as core evaluation indicators as comparison standards. See Table 3 for detailed results. The results show that our model shows superior performance over contrasting methods on most evaluation datasets and corresponding metrics, reaching an accuracy of 84.63% on the hybrid dataset we constructed. This clearly confirms the effectiveness and advantages of combining pre-trained models with efficient fine-tuning of parameters within a multi-task text framework. Although the optimal performance is not achieved on the specific street view text recognition task, we believe this is because the optimal parameter weights of the TrOCR architecture are not disclosed, and we can only adopt the non-optimal weight parameters of the TrOCR architecture, which means that we There is still space for improvement in our model performance.
Table 3: Different models were simultaneously evaluated in IAM,SROIE, STR benchmark and Mix Dataset average CER and average F1 score.
Model IAM SROIE STR BENCHMARKS MIX DATASETS
@@
CRNN(Shi,Bai,and Yao 2016)[26] 66.30 10.8 65.87 80.92 - 84.47 79.55 76.63
SRN(Yu et al.2020)[34] 77.36 9.11 72.12 87.33 - 88.66 82.71 81.16
VisionLAN(Wang et al.2021)[30] 75.68 7.99 77.65 92.22 - 90.07 81.69 83.11
ABINet(Fang et al.2021)[5] 78.66 8.87 79.88 90.75 - 89.22 83.47 84.19
@@
4.3 Ablation Study.
In order to prove the optimal combination of functions of the PEFT method, we designed a set of ablation experiments, and the result is shown in Table 5 and figure4. When encoder uses DoRA method and decoder uses LoRA method, the performance is significantly improved, with CER score of 5.42 and F1 score of 85.07, showing the best overall performance. The results of ablation experiments show that:
  • Full-parameter fine-tuning of only the image encoder is even better than fine-tuning of the image encoder and text decoder at the same time, because the pre-trained text decoder already has strong generalization in text decoding. performance, full-parameter fine-tuning on our relatively small mixed data set may destroy the original generalization.
  • Applying PEFT technology to encoders is far better than applying it to decoders. We believe the main reasons are:
    • 1.
      Diversity of information: Image encoders need to process diverse information from different types of images, while text decoders process relatively homogeneous text. Therefore, it may not be obvious whether the text decoder is processed or not, but it will be very important whether the image encoder is processed or not.
    • 2.
Table 5: Ablation experiments were performed in encoders and decoders using LoRA and DoRA methods and full parametric fine-tuning.

### v2

@@ Abstract
With the continuous development of OCR technology and the expansion of application fields, text recognition in complex scenes has become a key challenge. Factors such as multiple fonts, mixed scenes and complex layouts seriously affect the recognition accuracy of traditional OCR models. Although OCR models based on deep learning have performed well in specific fields or similar datasets in recent years, the generalization ability and robustness of the model are still a big challenge when facing complex environments with multiple scenes. Furthermore, training an OCR model from scratch or fine-tuning all parameters is very demanding on computing resources and inference time, which limits the flexibility of its application. This study focuses on a fundamental aspect of mixed text recognition in response to the challenges mentioned above, which involves effectively fine-tuning the pre-trained basic OCR model to demonstrate exceptional performance across various downstream tasks. To this end, we propose a parameter-efficient mixed text recognition method based on pre-trained OCR Transformer, namely DLoRA-TrOCR. This method embeds DoRA into the image encoder and LoRA into the internal structure of the text decoder, enabling efficient parameter fine-tuning for downstream tasks. Experimental results show that compared to similar parameter adjustment methods, our model DLoRA-TrOCR has the smallest number of parameters and performs better. It can achieve state-of-the-art performance on complex scene datasets involving simultaneous recognition of mixed handwritten, printed and street view texts.
@@ 1 Introduction
Figure 1: Optical Character Recognition: Text Detection and Text Recognition.
@@
Our research direction is to expand the application scope of pre-training fine-tuning paradigm in OCR. To construct a versatile multi-scene text recognition model, we built a dataset containing a mixture of handwritten, printed, and scene text. We then leveraged the pre-trained Transformer OCR model TrOCR, utilizing its weights from training on synthetic handwritten and printed text, and fine-tuned it on our mixed dataset. Introducing parameter-efficient fine-tuning (PEFT) techniques, we adopted the DoRA and LoRA methods for the encoder and decoder, respectively. This significantly reduced the trainable parameters while effectively improving performance on the mixed dataset. Experiments demonstrated outstanding performance in handling complex scenarios with mixed handwritten, printed, and scene text using only a small number of trainable parameters.
@@ 2 Related Work
Yu et al[34] propose the SRN, which utilizes a Transformer-encoder structure succeeded by a convolutional feature network to enhance visual features. Concerning the decoder, Yang et al[32] introduce the SRACN text detection algorithm that adopts a Transformer-decoder in lieu of the LSTM module, further substantiating the parallel training’s efficacy and precision benefits. Sheng et al[26] present the NRTR, which employs a complete Transformer architecture for both encoding and decoding input images. It only uses a few simple convolutional layers for deep feature extraction, verifying the effectiveness of the Transformer structure in text recognition. Lastly, Li et al[13] proposed the TrOCR model, applying the pre-trained Transformer encoder-decoder framework to OCR tasks. Distinctively, TrOCR eschews the CNN backbone typical for image comprehension, opting instead to directly utilize a pre-trained image Transformer model as the encoder and a pre-trained text Transformer model as the decoder.
@@
The foundation model[37]has strong generalization ability, and more and more research attempts to use transfer learning to apply it to tasks in specific scenarios. However, the scale of modern deep learning and the requirements for computing scale make traditional transfer learning strategies gradually ineffective, and the need for parameter-efficient adaptations become increasingly urgent[5]. Parameter-Efficient Fine-Tuning (PEFT)[20] aims to reduce the high cost of fine-tuning large models. This type of method achieves adaptation to downstream tasks by tuning a relatively small subset of parameters, rather than tuning all parameters of the model as in traditional methods. This strategy can significantly reduce computational cost and storage requirements compared to the total parameter amount.
Existing PEFT methods can be mainly divided into two categories: addition-based methods and reparameterization-based methods.Our research focuses on the second category of reparameterization-based methods, which are implemented by converting the adaptive parameters in the optimization process into a parametrically efficient form. This method was originally applied to natural language models. The representative of this type of method is Low-Rank Adaptation(LoRA)[9], which assumes that changes in weights during model tuning have low intrinsic rank. Based on this assumption, LoRA proposes a low-order decomposition optimization approximation method for the variation of the original weight matrix in the self-attention module. LoRA can match the fine-tuned performance of the GLUE benchmark while adding almost no additional inference overhead. Subsequently, Liu et al. introduced the idea of weight normalization based on LoRA, decomposed the pre-training weight into two components, amplitude and direction, and used LoRA to update the direction component. This method, called DoRA (Weight-Decomposed Low-Rank Adaptation)[33], enhances the learning capabilities and training stability of LoRA on some tasks without introducing any additional inference overhead.
Figure 2: Transformer-based architecture, where the encoder-decoder model is a pre-trained image transformer as the encoder and a pre-trained text transformer as the decoder.Then, this model is pre-trained in two stages on a synthetic dataset of millions of handwritten and printed texts.
@@ 3.1 Transformer-based OCR
Figure 3: Framework diagram of LoRA and DoRA methods.It is intuitive to see that after DoRA decomposes the pre-trained weights into directions and magnitudes, it updates the increments using LoRA in directions before computing with magnitudes.
@@ 3.2 Parameter-efficient Fine-tuning
where $B\in\mathbb{R}^{m\times r},A\in\mathbb{R}^{r\times n}$ are two small matrices, called LoRA weights, which are parameters that need to be learned during the fine-tuning process. $r$ is a hyperparameter that controls the rank of LoRA weight. When $r\ll\min(m,n)$, the number of parameters required by the LoRA weight will be much smaller than the original model weight matrix. By adding this low-rank update, LoRA enables efficient fine-tuning of pre-trained models. During the inference phase, the fine-tuned LoRA weights are combined with the original model weights to obtain the final inference weights. Since only two small matrices are stored, LoRA greatly reduces the additional parameter storage space required for fine-tuning. In our text decoder, we apply LoRA to both the query, key and value projection matrices of the self-attention module of the decoder and the encoder-decoder self-attention module, as well as the linear layer of the output. This parameter-efficient fine-tuning method allows us to significantly reduce the number of parameters required to fine-tune the text decoder while maintaining performance. Specifically, the low-rank matrix of LoRA is used to reconstruct the multi-head attention mechanism in the Transformer block, and its incremental update can be expressed as
$\operatorname{head}_{i}=\operatorname{self-att}({Q}_{i},{K}_{i},{V}_{i})$
$Q_{i}=(W_{q,i}+B_{q,i}A_{q,i})\mathbf{x}$
$K_{i}=(W_{k,i}+B_{k,i}A_{k,i})\mathbf{x}$
$V_{i}=(W_{v,i}+B_{v,i}A_{v,i})\mathbf{x}.$
Each attention head has its own query, key, value matrix and its corresponding LoRA low-rank matrix incremental update term ($B_{*,i}A_{*,i}$). After splicing the results of different attention heads, they are fused by a weight matrix, and also updated by LoRA low-rank matrix:
$O=(W_{o}+B_{o}A_{o})\operatorname{Concat}(\operatorname{head}_{1};\cdots;\operatorname{head}_{h})\$
@@
$Q_{i}$ $=(W_{q,i}+\|W_{q,i}\|_{c}\frac{W_{q,i}+B_{q,i}A_{q,i}}{\|W_{q,i}+B_{q,i}A_{q,i}\|_{c}})\mathbf{x}$
$K_{i}$ $=(W_{k,i}+\|W_{k,i}\|_{c}\frac{W_{k,i}+B_{k,i}A_{k,i}}{\|W_{k,i}+B_{k,i}A_{k,i}\|_{c}})\mathbf{x}$
$V_{i}$ $=(W_{v,i}+\|W_{v,i}\|_{c}\frac{W_{v,i}+B_{v,i}A_{v,i}}{\|W_{v,i}+B_{v,i}A_{v,i}\|_{c}})\mathbf{x}.$
@@
Table 1: Optimization of computational resources and model trainable parameters using LoRA, DoRA and DLoRA methods, which load the pre-trained weights of the model using fp16.
Model PEFT Method GPU Memory Tranable Params Total Params # Params
@@
Table1 details the computing resource consumption and parameter scale of TrOCR-base model under different PEFT strategies. Specifically, the following situations are compared in the table:
  • 1. Full LoRA or DoRA: the encoder and decoder both adopt LoRA or DoRA method;
  • 2. Hybrid configuration: the encoder uses DoRA method, while the decoder adopts LoRA method;
  • 3. Full parameter fine-tuning: no PEFT method, all parameters participate in the training.
For the model using PEFT method, half-precision (fp16) is used to load pre-trained weights and freeze. This strategy only accounts for 0.6% of the trainable parameters, and the amount of GPU memory is reduced by 30% under the same batch size. In fact, the precision of the floating-point numbers can be reduced even further, thus significantly increasing the batch size and improving the training efficiency while maintaining the same amount of memory.
@@ 4.1 Experiment Setup
From the IAM dataset[21], which contains handwritten English text and is one of the most commonly used datasets in the field of handwritten text recognition. We use the partitioning of the Aachen dataset: the training set contains 6,482 lines of text in 747 forms, the validation dataset contains 966 lines of text in 115 forms, and the test dataset contains 2,915 lines of text in 336 forms.
@@
From the task2 of SROIE dataset[10], which focuses on text recognition in receipt images. The SROIE training dataset and test dataset contain 626 and 361 receipt images respectively. After cropping and removing duplicate images, the training dataset contains 10,682 lines of text, and the test dataset contains 6,897 lines of text.
@@
From multiple public benchmarks, including IIIT5K-Words (IIIT5K)[18], Street View Text(SVT)[30], ICDAR 2013 (IC13)[11], ICDAR 2015(IC15)[12], Street View Text-Perspective (SVTP)[19], and CUTE80 (CUTE)[25]. These datasets contain various scene text images with noise such as blur, occlusion, or low resolution. The training set contains 7573 lines of text, and the test set contains 11435 lines of text. The specific dataset situation is shown in the Table2.
@@
To achieve better generalization on multi-scene tasks, we adopt hybrid measures on the dataset. Compared with the circular sampling method [28]that may lead to errors in uneven data distribution. We adopt a direct and balanced data integration strategy, directly mixing the datasets of the above handwritten , printed and scene text. Such a mixed dataset eliminates the possibility of a single task dominating the training process, and promotes the comprehensive learning of the model for a variety of image types and corresponding text descriptions. All subsequent comparison, combination and ablation experiments are trained and tested based on this mixed dataset.
@@
We considered three evaluation metrics that are widely used in OCR systems, Character Error Rate(CER) , F1 score and Word Accuracy. CER is designed to be evaluated at the character level, which is case-sensitive, while the F1 score and word accuracy are evaluated at the word level, which are case-insensitive. CER is defined as the edit distance between the predicted text and the ground truth. CER can be calculated as:
@@
WER and CER are similar benchmark scores. The main difference is that the smallest unit of text in the calculation method is changed from characters to words, so that WAR can be calculated by $1-\mathrm{WER}$. We mainly use it in the evaluation of the STR benchmark datasets.
@@
In the model configuration, we use the official pre-training weights of the TrOCR-Base model, and use the AdamW optimizer to perform training for 20 epochs and a batch size of 16. The initial learning rate of full-parameter Fine-Tuning is 1e-5, and the initial learning rate of PEFT method is 5e-5. Our comparative model experiments are all based on PyTorch , and all experiments are performed on an Nvidia RTX 4090 GPU with 24 GB of memory.
In order to accurately evaluate the model performance and ensure the rationality of the training, validation and testing processes, we divided the mixed dataset into a training dataset and a validation dataset according to the classic 9:1 ratio. In addition, we also retain an independent test dataset before mixing to ensure that the evaluation of the final performance of the model is completely independent of the training process.
@@ 4.2 Experiment Result
Table 3: Different models were simultaneously evaluated in IAM,SROIE, STR benchmark and Mixed Dataset average CER and average F1 score.
Model IAM SROIE STR BENCHMARKS MIXED DATASETS
@@
CRNN(Shi,Bai,and Yao 2016)[27] 66.30 10.8 65.87 80.92 5.54 84.47 79.55 76.63
CLOVA OCR(Baek et al. 2019)[2] 67.71 9.24 70.77 84.41 4.88 86.23 79.55 76.63
SRN(Yu et al.2020)[35] 77.36 9.11 72.12 87.33 4.91 88.66 82.71 81.16
VisionLAN(Wang et al.2021)[31] 75.68 7.99 77.65 92.22 2.74 90.07 81.69 83.11
ABINet(Fang et al.2021)[6] 78.66 8.87 79.88 90.75 3.12 89.22 83.47 84.19
@@
The evaluation covers the above-mentioned benchmark datasets, and uses F1 and CER as core evaluation indicators as comparison standards. The results show that our model has superior performance over contrasting methods on most evaluation datasets and corresponding metrics, reaching an accuracy of 84.63% on the mixed dataset we constructed. This clearly confirms the effectiveness and advantages of combining pre-trained models with PEFT within the multi-task OCR framework. Although the optimal performance is not achieved on the specific street view text recognition task, we believe this is because the optimal parameter weights of the TrOCR architecture are not disclosed, and we can only adopt the non-optimal weight parameters of the TrOCR architecture, which means that we There is still space for improvement in our model performance.
@@
Table 5: Comparative results of ablation experiments in encoder and decoder using LoRA and DoRA methods and full parameter fine-tuning.
@@
Figure 4: The improvement of DoRA and LoRA methods on the encoder and decoder modules respectively on the performance of our model.
4.3 Ablation Study.
In order to prove the optimal combination of functions of the PEFT method, we designed a set of ablation experiments, and the result is shown in Table5 and figure4. When encoder uses DoRA method and decoder uses LoRA method, the performance is significantly improved, with CER score of 5.42 and F1 score of 85.07, showing the best overall performance.
The results of ablation experiments show that:
  • Only fine-tuning the image encoder with full parameters is even better than fine-tuning the image encoder and the text decoder together. This is because the pre-trained text decoder already has strong generalization in text decoder, and fine-tuning it with full parameters on our relatively small mixed dataset would destroy the generalization.
  • Applying PEFT technology to encoders is far better than applying it to decoders. Figure 4 shows the difference in average CER and average F1 using LoRA, DoRA and full parameter fine-tuning for encoder and decoder respectively. We believe the main reasons are:
    • 1.
      Diversity of information: Image encoders need to process diverse information from different types of images, while text decoders process relatively homogeneous text. Therefore, it may not be obvious whether the text decoder is processed or not, but it will be very important whether the image encoder is processed or not.This shows that the image encoder is more task-related in OCR.
    • 2.